# Multiple Stochastic Prompt Tuning for Practical Cross-Domain Few Shot Learning

### v1

@@ 6 Conclusion
In this work, we advocate a more realistic and challenging setting termed as the practical CDFSL (pCDFSL) setting, where a large-scale model like CLIP is adapted to a few-shot downstream dataset without the episodic constraints. We further propose a novel prompt tuning approach, MIST to address the various challenges of this setting. Different from existing prompt tuning approaches, our proposed MIST learns multiple text prompts per class, modeled by distinct learnable Gaussian distributions to represent the inherent multimodal class distributions as well as mitigate overfitting. Extensive experiments on the proposed pCDFSL setting as well as additional analysis show the effectiveness of our proposed approach compared to state-of-the-art methods.
Supplementary Material
Aircraft DTD Flowers Caltech101 Food101 Oxford Pets Average
MaPLe 29.03 54.73 80.80 94.30 86.90 92.05 72.97
PromptSRC 32.80 60.64 91.31 94.77 86.06 93.23 76.47
TCP 36.20 63.97 94.40 95.00 85.30 91.90 77.80
MIST 36.60 61.40 91.73 94.67 82.47 92.80 76.61
Table 6: Performance (%) on the CLIP-based FSL datasets for the 4-shots setting. Highest and second highest values are marked by bold and underline respectively.
A. More Dataset details: We have used the four datasets from the BSCDFSL [10] benchmark to construct our proposed pCDFSL setting. This benchmark consists of images and classnames from various specialized domains like medical, satellite and agricultural fields, displaying extreme domain shifts from natural images. We provide more details on each of these datasets below:
1) EuroSAT [11]: This dataset consists of satellite images of various terrains and comprises of 10 classes, namely, Annual Crop, Forest, Herbaceous Vegetation, Permanent Crop, Residential Buildings, Pasture, Industrial buildings, Highway, River, Sea-Lake.
2) ISIC [5]: Consists of dermascopic skin disease images and has 7 classes, namely, Melanoma, Melanocytic Nevi, Basal Cell Carcinoma, Actinic Keratosis, Benign Keratosis, Dermatofibroma, Vascular Lesions.
3) Plant Disease [22]: Contains images of leaf diseases across 38 classes, e.g., Apple Scab, Apple Black Rot, Apple Cedar Rust, Apple Healthy, Blueberry healthy, Cherry Powdery Mildew, Cherry Healthy, and so on.
4) ChestX [27]: This dataset comprises of greyscale chest X-Ray images across 7 classes, namely, Atelectasis, Cardiomegaly, Effusion, Infiltration, Mass, Nodule, Pneumothorax.
The datasets ranked in decreasing order of similarity with ImageNet [6] are as follows: Plant Disease $>$ EuroSAT $>$ ISIC $>$ ChestX.
B. CLIP-based FSL performance: The detailed results of the few-shot experiments on the CLIP-based FSL datasets are reported in Table 6.
C. Memory consumption: We report the comparative GPU memory consumptions for the 1-shot case in Table 7. Here, we observe that our proposed MIST utilizes slightly more memory than MaPLe [15] during training, due to addition of learnable parameters, but utilizes the same memory during inference. However, PromptSRC [16] takes up much more memory during both training and inference. TCP [29] takes slightly less memory due to fewer learnable parameters during training, but consumes a huge memory during inference, due to loading of the frozen CLIP model. Overall, our proposed MIST framework fairs comparably to MaPLe, and is more suitable for real-world deployment compared to the other methods.
MaPLe PromptSRC TCP MIST (Ours)
Training 1172 2168 778 1348
Inference 1608 2994 4946 1610
Table 7: Comparison of the GPU memory requirements (in MBs) for training and inference.

### v2

@@ 6 Conclusion
In this work, we propose a novel framework, MIST for adapting foundation VLMs like CLIP to realistic few-shot scenarios characterized by extreme domain and label semantic shifts. Motivated by the limitations of existing parameter efficient fine-tuning approaches, we incorporate multiple text prompts per class, modeled by distinct learnable Gaussian distributions to represent the inherent multimodal class distributions as well as mitigate overfitting. Extensive experiments on multiple benchmarks as well as additional analysis show the effectiveness of our proposed approach compared to state-of-the-art methods.